The hypotenuse of a right triangle is constantly 8 inches long. One leg is increasing at the

rate of 3in./min. At what rate is the area of the triangle changing when the increasing leg is

4in.


1
Expert's answer
2022-05-31T11:06:17-0400


Given

"x^2+y^2=8^2, x>0, y>0"

Then


"x=\sqrt{64-y^2}, 0<y<8"

The area of the triangle is


"A=\dfrac{1}{2}xy"

Substitute


"A=A(t)=\dfrac{1}{2}y\sqrt{64-y^2}"

Differentiate with respect to "t"


"\dfrac{dA}{dt}=\dfrac{1}{2}(\sqrt{64-y^2}+\dfrac{-2y^2}{2\sqrt{64-y^2}})\dfrac{dy}{dt}"

"\dfrac{dA}{dt}=(\dfrac{64-y^2-y^2}{2\sqrt{64-y^2}})\dfrac{dy}{dt}"

"\dfrac{dA}{dt}=(\dfrac{32-y^2}{\sqrt{64-y^2}})\dfrac{dy}{dt}"

If "\dfrac{dy}{dt}=3\ in/min, y=4\ in," then


"\dfrac{dA}{dt}=(\dfrac{32-4^2}{\sqrt{64-4^2}})(3)=4\sqrt{3}({in}^2/min)"

The area of the triangle  is increasing at the rate of "4\sqrt{3}" in2/min.